# Di-NeRF: Distributed NeRF for Collaborative Learning with Relative Pose Refinement

### v2

@@ II Literature Review
Extending NeRF to multi-robot scenarios has also been proposed in Block-NeRF [8], NeRFuser [29], and NeRFusion [30]. These methods use blending techniques to render an image, often done centralized, performing inference on multiple NeRF models. Furthermore, it is assumed that there is prior knowledge about the relative pose of the robots. A similar research is federated learning [31, 19], where the training is done in a decentralized optimization, and each node learns a common NeRF in parallel. Then the weights are transferred to a server for aggregation. In DiNNO [10], a distributed algorithm for collaborative learning of a neural model is presented, demonstrating a range of experiments including multi-agent RL, MNIST classification, and 2D neural implicit mapping with known relative poses. The core of the algorithm is a distributed optimization method known as C-ADMM  [1].
@@ III-A Problem Statement
Assume $N$ robots connected to each other with a communication graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ consisting of vertices $\mathcal{V}=\{1,...,N\}$ and edges $\mathcal{E}$, on which pairwise communication can occur. Given a set of images $\boldsymbol{I}_{i}$ captured by robot $i$, $i=1,...,N$, from a scene with the associated local camera poses, $\boldsymbol{\pi}_{i}$, the goal of Di-NeRF is to build a scene representation, $\theta$, that enables all the robots to generate realistic images from novel viewpoints. Here, NeRF [18] is employed as the model for 3D reconstruction, $\theta$. Since the poses of each robot are defined in a local coordinate system, it is necessary to perform relative pose optimization to build a consistent model that accurately represents the entire scene. The problem is expressed as:
@@
$\boldsymbol{d}=R\begin{pmatrix}\frac{u-W/2}{f}&\frac{{\color[rgb]{0,0,0}-}v{\color[rgb]{0,0,0}+}H/2}{f}&-1\end{pmatrix}^{\top}\leavevmode\nobreak\ ,$ (2)
@@ III-B Relative Camera Pose Optimization
In multi-robot NeRF, robots do not send raw data over the wireless network. Therefore, using an SfM pipeline to calculate the relative poses for all robots is not easy. Here, the camera poses of each robot are expressed based on a local coordinate system, and gradually all poses transfer to the robot $R^{g}$ coordinate system (throughout this paper, the robot $R^{g}$ will be the arbitrary reference agent.). The optimization of the relative poses is performed jointly with the model parameters. Relative camera poses of multi-robot systems can be expressed as a camera-to-world transformation matrix (similar to each robot’s local camera pose) $T_{i}^{g}=[R^{g}|t^{g}]$in $SE(3)$, where $R^{g}\in SO(3)$ and $t^{g}\in\mathbb{R}^{3}$ show camera rotation and translation, respectively [15]. Optimizing the translation vector $t^{g}$ involves the designation of trainable parameters, given its definition in Euclidean space. For camera rotation, which is in $SO(3)$ space, the axis-angle representation is chosen: $\phi:=\alpha\omega,\phi\in\mathbb{R}^{3}$, where $\omega$ and $\alpha$ are a normalized rotation axis and a rotation angle, respectively. The rotation matrix $R^{g}$ can be expressed from Rodrigues’ formula:
@@
Training NeRF and the relative camera poses jointly adds another degree of ambiguity to the problem, and causes the NeRF optimization to converge to local minima, especially for large relative poses. To avoid local minima, we use monocular depth estimation [17]. However, the depth maps suffer scale and shift distortions across frames, making mono-depth maps not multi-view consistent. To utilize mono-depth effectively, we explicitly optimize scale and shift parameters during NeRF training by penalizing the difference between rendered depth and mono-depth. For each image $I^{m}_{i}$ (out of $M_{i}$ images) of robot $i$, we generate mono-depth sequence $D^{m}_{i},m=0,...,M_{i}-1$ from input images with an off-the-shelf monocular depth network, i.e., DPT [32]. Two linear transformation parameters $\boldsymbol{\psi}_{i}^{m}=(\alpha^{m}_{i},\beta^{m}_{i}),m=0\ldots M_{i}-1$ for each mono-depth map as scale and shift factors can recover a multi-view consistent depth map ${D^{m}_{i}}^{*}=\alpha^{m}_{i}D^{m}_{i}+\beta^{m}_{i}$, for each image $m$ in each robot $i$. This joint optimization of distortion parameters and NeRF is achieved by adding a depth loss $L^{depth}_{i}=\sum_{m}||{D^{m}_{i}}^{*}-\hat{D}^{m}_{i}||$ to the RGB loss of NeRF [18], $L^{RGB}_{i}=\sum_{m}||{I^{m}_{i}}-\hat{I}^{m}_{i}||$. Here $\hat{D}^{m}_{i}$ and $\hat{I}^{m}_{i}$ are NeRF-rendered depth/RGB images, described in [17][18]. Thus, the local NeRF loss for each robot $i$ is defined as:
@@ III-C Distributed Formulation
[…]
@@ III-D Distributed Optimization of NeRF
Alg. 1 summarizes the optimization procedures for local primal, dual variables, and relative poses for each robot. In line 1, The initial values for the NeRF parameters $\theta_{i}^{(0)}$, the dual variable $y_{i}^{(0)}$, and the depth distortion parameters $\boldsymbol{\psi}_{i}^{(0)}$ are set to zero. The initial value for the relative poses - three rotation angles $R$ and three translation displacements $t$ are explained in Sec. IV. The first robot $R^{g}$ undergoes the NeRF training process (ln 6-7), communicates the learned weights with other robots (ln 10-11), and updates the dual variable (ln 12) for 200 steps. The remaining robots engage in collaborative optimization, simultaneously refining NeRF weights and adjusting relative camera poses and depth distortion parameters in a distributed training framework based on the weights of robot $R^{g}$ (lines 8-9 for NeRF training, lines 10-12 for communication and dual update). This iterative communication and training process persists until uniformity is achieved in both the weights and optimized relative camera poses across all robots. The relative pose between robot $i$’s coordinate system and robot $R^{g}$’s coordinate system $T_{i}^{g}\in SE(3)$ is optimized jointly in the distributed training (ln 9).
@@
8:           $+\rho\sum_{j\in\mathcal{N}_{i}}\left\|\theta_{i}-\bar{\theta}^{(k)}\right\|_{2}^{2}\}$
@@
[…]
@@ IV-A Di-NeRF vs Baseline NeRF on Synthetic Dataset
Table I: Image and relative camera pose metric performances for two robots. PSNR and SSIM are reported for each robot. The relative pose accuracy is shown for $R^{2}$. $R^{1}$ is the global frame. lower/higher better is indicated by $\downarrow$/$\uparrow$ respectively.
@@ IV-B Di-NeRF vs Baseline NeRF on Real-world Dataset
#Robots PSNR $\uparrow$ SSIM $\uparrow$ $\delta R^{\circ}$downarrow $\delta t(cm)$downarrow
@@
Table III: Mean validation result of Di-NeRF for 5 robots, with different connectivities, for the Barn sequence. Di-NeRF works well even if the network of robots is not fully connected. lower/higher better is indicated by $\downarrow$/$\uparrow$ respectively.
@@ IV-C Trajectory Overlap Analysis
In this experiment, the sensitivity of the convergence of the relative poses with respect to the overlaps between the views of the robots is analyzed. For this experiment, two robots are considered. There are no common images between the robots, and the overlaps are only in the trajectories. Fig. 9 shows the top-down view of the overlaps, ranging from 5% to 40%, i.e. [5, 10, 20, 30, 40]%. The overlaps are determined using $x$ and $y$ coordinates of camera poses, assuming the poses are known in a coordinate frame. To divide the views among $N$ robots, first, the $x-y$ plane containing the desired object (the Chair, in this case) is divided into $N$ sectors, centered on the object. This will generate a 0% overlap for $N$ robots, which then is extended to increase the overlap, as shown in Fig. 9.
@@ IV-H Technical Details of the Experiments
In this paper, three datasets are used to evaluate Di-NeRF: the synthetic dataset [2] (800$\times$800 pixels), the Tanks and Temples [37] (1920$\times$1080 pixels), and the Waymo dataset [8] (1217$\times$1096 pixels). For all datasets, the batch size is 2048, the learning rate is 0.01, the weight for depth loss $\gamma$ in Eq. (4) is 0.01 and the value of $\rho$ in Eq. (9) is 0.001. $\rho$ is the weight for the quadratic term and is the step size in the gradient ascent for the dual variable optimization. The number of steps before a communication round is 200. There are 5 communication rounds in all experiments.
@@
Table IV: Di-NeRF and baseline (Bsln.) performances for the Mission Bay sequence for 6 robots. $R\#1$ is the reference frame.
@@
$\delta R^{\circ}\downarrow$ 2.89 2.86 2.81 2.91 2.79 -
$\delta t$(cm)$\downarrow$ - 4.2 4.0 4.5 4.2 4.7 -

### v3

@@ II Literature Review
Extending NeRF to multi-robot scenarios has also been proposed in Block-NeRF [8], NeRFuser [29], and NeRFusion [30]. These methods use blending techniques to render an image, often done centralized, performing inference on multiple NeRF models. Furthermore, it is assumed that there is prior knowledge about the relative pose of the robots. A similar research is federated learning [31, 19], where the training is done in a decentralized optimization, and each node learns a common NeRF in parallel. Then the weights are transferred to a server for aggregation. A memory-efficient multi-robot SLAM approach based on multi-block ADMM is proposed in [32]. In DiNNO [10], a distributed algorithm for collaborative learning of a neural model is presented, demonstrating a range of experiments including multi-agent RL, MNIST classification, and 2D neural implicit mapping with known relative poses. The core of the algorithm is a distributed optimization method known as C-ADMM [1]. To establish a shared reference frame, [33] proposes a method for incremental inference from arbitrary poses. Additionally, im2nerf [34, 16] enables neural radiance field prediction from a single image, addressing view sparsity.
@@ III-A Problem Statement
Assume $N$ robots connected to each other with a communication graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ consisting of vertices $\mathcal{V}=\{1,...,N\}$ and edges $\mathcal{E}$, on which pairwise communication can occur. Given a set of images $\boldsymbol{I}_{i}$ captured by robot $i$, i=1,,Ni=1,...,N, from a scene with the associated local camera poses, $\boldsymbol{\pi}_{i}$, the goal of Di-NeRF is to build a scene representation, $\theta$, that enables all the robots to generate realistic images from novel viewpoints. Here, NeRF [18] is employed as the model for 3D reconstruction, $\theta$. Since the poses of each robot are defined in a local coordinate system, it is necessary to perform relative pose optimization to build a consistent model that accurately represents the entire scene. The problem is expressed as:
@@
$\boldsymbol{d}=R\begin{pmatrix}\frac{u-W/2}{f}&\frac{{\color[rgb]{0,0,0}-}v{\color[rgb]{0,0,0}+}H/2}{f}&-1\end{pmatrix}^{\top}\leavevmode\nobreak\$, (2)
@@ III-B Relative Camera Pose Optimization
In multi-robot NeRF, robots do not send raw data over the wireless network. Therefore, using an SfM pipeline to calculate the relative poses for all robots is not easy. Here, the camera poses of each robot are expressed based on a local coordinate system, and gradually all poses transfer to the robot $R^{g}$ coordinate system (throughout this paper, the robot $R^{g}$ will be the arbitrary reference agent.). The optimization of the relative poses is performed jointly with the model parameters. Relative camera poses of multi-robot systems can be expressed as a camera-to-world transformation matrix (similar to each robot’s local camera pose) $T_{i}^{g}=[R^{g}|t^{g}]$ $\in$ $SE(3)$, where $R^{g}\in SO(3)$ and $t^{g}\in\mathbb{R}^{3}$ show camera rotation and translation, respectively [15]. Optimizing the translation vector $t^{g}$ involves the designation of trainable parameters, given its definition in Euclidean space. For camera rotation, which is in $SO(3)$ space, the axis-angle representation is chosen: $\phi:=\alpha\omega,\phi\in\mathbb{R}^{3}$, where $\omega$ and $\alpha$ are a normalized rotation axis and a rotation angle, respectively. The rotation matrix $R^{g}$ can be expressed from Rodrigues’ formula:
@@
[…]
@@ III-C Distributed Formulation
$\text{s.t.}\quad\theta_{i}$ $=z_{ij},\quad\forall j\in\mathcal{N}_{i}$
@@ III-D Distributed Optimization of NeRF
Alg. 1 summarizes the optimization procedures for local primal, dual variables, and relative poses for each robot. In line 1, The initial values for the NeRF parameters $\theta_{i}^{(0)}$, the dual variable $y_{i}^{(0)}$, and the depth distortion parameters $\boldsymbol{\psi}_{i}^{(0)}$ are set to zero. The initial value for the relative poses - three rotation angles $R$ and three translation displacements $t$ - are explained in Sec. IV. The first robot $R^{g}$ undergoes the NeRF training process (ln 6-7), communicates the learned weights with other robots (ln 10-11), and updates the dual variable (ln 12) for 200 steps. The remaining robots engage in collaborative optimization, simultaneously refining NeRF weights and adjusting relative camera poses and depth distortion parameters in a distributed training framework based on the weights of robot $R^{g}$ (lines 8-9 for NeRF training, lines 10-12 for communication and dual update). This iterative communication and training process persists until uniformity is achieved in both the weights and optimized relative camera poses across all robots. The relative pose between robot $i$’s coordinate system and robot $R^{g}$’s coordinate system $T_{i}^{g}\in SE(3)$ is optimized jointly in the distributed training (ln 9).
@@
8:           $\rho\sum_{j\in\mathcal{N}_{i}}\left\|\theta_{i}-\bar{\theta}^{(k)}\right\|_{2}^{2}\}$
@@
10:     $Γ{\color[rgb]{0,0,0}{\Gamma}}_{i}^{(k+1)}=\underset{{\color[rgb]{0,0,0}{\Gamma}}}{\operatorname{argmin}}$
@@ IV-A Di-NeRF vs Baseline NeRF on Synthetic Dataset
Table I: Image and relative camera pose metric performances for two robots. PSNR and SSIM are reported for each robot. The relative pose accuracy is shown for $R^{2}$. $R^{1}$ is the global frame. lower/higher better is indicated by $\downarrow$$\uparrow$ respectively.
@@ IV-B Di-NeRF vs Baseline NeRF on Real-world Dataset
#Robots PSNR $\uparrow$ SSIM $\uparrow$ $\delta R^{\circ}$ $\downarrow$ $\delta t(cm)$ $\downarrow$
@@
Table III: Mean validation result of Di-NeRF for 5 robots, with different connectivities, for the Barn sequence. Di-NeRF works well even if the network of robots is not fully connected. lower/higher better is indicated by $\downarrow$$\uparrow$ respectively.
@@ IV-C Trajectory Overlap Analysis
In this experiment, the sensitivity of the convergence of the relative poses with respect to the overlaps between the views of the robots is analyzed. For this experiment, two robots are considered. There are no common images between the robots, and the overlaps are only in the trajectories. Fig. 9 shows the top-down view of the overlaps, ranging from 5% to 40%, i.e. [5, 10, 20, 30, 40]%. The overlaps are determined using $x$ and $y$ coordinates of camera poses, assuming the poses are known in a coordinate frame. To divide the views among $N$ robots, first, the xyx-y plane containing the desired object (the Chair, in this case) is divided into $N$ sectors, centered on the object. This will generate a 0% overlap for $N$ robots, which then is extended to increase the overlap, as shown in Fig. 9.
@@ IV-H Technical Details of the Experiments
[…]
@@
Table IV: Di-NeRF and baseline (Bsln.) performances for the Mission Bay sequence for 6 robots. R#1R\#1 is the reference frame.
@@
$\delta R^{\circ}\downarrow$ - 2.89 2.86 2.81 2.91 2.79 -
[…]